how do you do this?
Write the slope intercept form of the equation of the line passing through the point (5,-4) and perpendicular to the line y=-4/3x+5

Answer by mananth(16946)

1 y = -1 1/ 3 x + 5
Divide by 1
y = -4/ 3 x + 5
Compare this equation with y=mx+b, m= slope & b= y intercept
slope m = -4/3

The slope of a line perpendicular to the above line will be the negative reciprocal 3/4
m1*m2=-1
The slope of the required line will be 3/4

m= 3/4 ,point ( 5 , -4 )
Find b by plugging the values of m & the point in
y=mx+b
-4 = 15/4 + b
b= -31/4
m= 3/4
The required equation isy = 3/ 4x- 31/ 4
m.ananth@hotmail.ca